In any triangle, A=62 degrees, c=15, and b=12; find the area.
Answer by sarah_adam(201)

Side Angle Side :
when length of two sides and one angle is given then the formula for calculating area is A = (1/2 *a*b* sin theta )where a and b are the sides of the triangle and theta is the angle included between the two sides
here sides are 15 and 12 and theta is 62 degrees
A = 1/2 * 15 *12 *sin(62)
A = 80 sin (62)
A = 80 * 0.88 (sin 62 = 0.88)
A = 70.63 sq units
(please verify the value of Sin 62)